Evaluate the surface integral. zds Sis the surface x = y + 322, 0 sy s 1,0 S 234
Answer #1

#Given SS 2ds Sto where x = y + 32², 05461, 06224 Sol => SS2ds Where ds = ſit (d2 2 + ( didz x= y + 322 x29+ 32² Given dx dz